Find the minimum magnetic field needed for the Zeeman effect to be observed


in a spectral line of 400 nm wavelength when a spectrometer whose resolution


is 0.01nm is used.

1
Expert's answer
2021-12-15T22:33:19-0500

We know that

"\nu=\frac{c}{\lambda}"

"\nu=\frac{3\times10^8}{400\times10^{-9}}=7.5\times10^{-14}Hz"

Again

"\nu_1=\frac{c}{\lambda_1}"

"\nu_1=\frac{3\times10^8}{400\times10^{-9}+0.010\times10^{-9}}"

"\nu_1=\frac{3\times10^8}{400\times10^{-9}}\times\frac{1}{1+\frac{0.1}{400}}"

"∆\nu=7.5\times10^{14}(1-2.5\times10^{-5})"

"∆\nu=1.875\times10^{10}Hz"

Magnetic field

"B=\frac{4\pi m∆\nu}{e}"


"B=\frac{4\times3.14\times9.1\times10^{-31}\times1.875\times10^{10}}{1.06\times10^{-19}}=1.34T"